OpenStudy (anonymous):

20 percent out of 80 is what out of a 100

OpenStudy (jdoe0001):

let's say 80 is the 100%, what is the quantity 20 out of 80 in percent? well \(\large \begin{array}{ccllll} quantity&percentage\\ \hline\\ 80&100\\ 20&x \end{array}\implies \cfrac{80}{20}=\cfrac{100}{x}\) once you get "x", then do the same to find what is "x % of 100"

OpenStudy (mathstudent55):

The way you stated your problem, this is how you do it. 20% of 80 = = 20/100 * 80 = 0.2 * 80 = 16 16 out of 100 is 16% since % means out of 100. Perhaps, you didn't mean it the way you stated it.